OpenStudy (anonymous):

hwy does anyone know the derivative of f(x)=3x+5x^-1 ?

OpenStudy (saifoo.khan):

derivative of 3x = x derivative of 5x^-1 = -5x^-2

OpenStudy (saifoo.khan):

So collectively, \[\Large 3x+ 5x^{-1} \to -\frac{5}{x^2} +x\]

OpenStudy (anonymous):

Holy cow. I never coulda figgered that on my own.

OpenStudy (saifoo.khan):

lol @ holycow

OpenStudy (anonymous):

I get the one with the negative exponent. But doesn't the 3x derive into just 3? (1 x 3=3, and the x gets dropped...)?

OpenStudy (saifoo.khan):

Oh man! my mistake! holy cow!

OpenStudy (anonymous):

that is what wass thing pinhead I was thinking, 3 - 5x^-2

OpenStudy (saifoo.khan):

i mean it's 3 not x.

OpenStudy (anonymous):

haha!

OpenStudy (saifoo.khan):

Sowwie.

OpenStudy (anonymous):

soooo, f'(x)= 3-5x^-2 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks so much, you two. I am so very bad at math. I can do other subjects really well, but math is holding me back

OpenStudy (anonymous):

welcome to the website then - it may be a great resource

OpenStudy (anonymous):

It sure seems like it! Another student friend o mine recommended it

OpenStudy (anonymous):

now how can I factor this?

OpenStudy (anonymous):

new question? if so do new post

OpenStudy (anonymous):

I am trying to find critical numbers so i can find relative mins/max

OpenStudy (anonymous):

f'(x)= 3-5x_2....I need to factor this, tosolve for 0, so I can find critical numbers...

OpenStudy (anonymous):

f'(x)=3-5x^-2

OpenStudy (anonymous):

how can I factor f'(x)= 3-5x^-2 ?

OpenStudy (anonymous):

there's no way! How can I solve for X and find my critical values?

OpenStudy (anonymous):

 look at it this way- does that hel you